# Metacognition and Uncertainty Communication in Humans and Large Language Models

### v1

@@ 4.2 Divergences Between Human and LLM Metacognition
Despite a number of parallels, there remain important differences between human and LLM metacognition. In humans, the ability to metacognitively form confidence judgments is typically thought to rest on the formation of a second-order representation: a separate evaluation or reassessment of the internal representations prompted by input information and which gave rise to a behavioral output (Peters, 2022); unless explicitly present in their architecture, LLMs may not form such second-order self-evaluative representations unless explicitly prompted to do so. Relatedly, LLMs may be less able to correctly evaluate the source of such second-order representations or judgments, suggesting they lag humans in distinguishing between metacognition and theory of mind. LLMs are prone to conflate their own beliefs with those attributed to others; that is, they are less able to separate the speaker’s belief from their own compared to humans when interpreting uncertain statements (Belém et al., 2024).
Another difference may stem from the domain generality or specificity of metacognition in humans. It is thought that some shared processes underlie metacognition about perception, memory, and cognition may exist and rely on common neural structures, while others may be domain specific—i.e., separable computational or neural modules for perceptual versus cognitive or memory metacognition (Morales et al., 2018). As LLMs are increasingly integrated into many highly different tasks and reasoning domains, attending to their domain-specific versus domain-general metacognitive capacities will become increasingly urgent (see, e.g., (Griot et al., 2025) for LLMs’ metacognitive failures in medical reasoning).
@@ 5 Communication of Uncertainty in Human-AI Interaction
As discussed above, metacognitive sensitivity is the degree to which confidence judgments can discriminate between right and wrong answers, which is critical to effective decision-making in humans (Fleming, 2024). For optimal interaction and humans’ trust of AI systems, LLMs thus must be able to convey to human deciders not only whether their decisions are likely to be correct (Lee et al., 2024; Steyvers et al., 2025; Li & Steyvers, 2025), but also whether they are likely to respond capably to novel tasks (Kadavath et al., 2022). Problematically, LLMs appear reluctant to express uncertainty (Zhou et al., 2024). Because humans rely heavily on linguistic uncertainty expressions (Steyvers et al., 2025; Zhou et al., 2024), the absence of expressions of uncertainty may raise humans’ reliance on model outputs even beyond the already-overconfident judgments the models express. A potential reason for LLMs’ lack of uncertainty expression may stem from training protocols, which fine-tune LLM outputs stylistically to align with human preferences for such responses—which themselves present biases against responses that contain uncertainty (Zhou et al., 2024). Unfortunately, this problem may be further exacerbated as LLMs are used for increasingly challenging applications, potentially by increasingly non-expert users. Because individuals who do not possess topical expertise are less able to correctly assess the expertise of others (Bower et al., 2024), non-expert individuals working with LLMs on complex problems might be especially susceptible to influence by superficial features of their response, like the lack of uncertainty expressions or even the length of the answer (Steyvers et al., 2025). Humans and LLMs may also rely on different sets of cues when assessing their confidence in other humans, such as humans’ reliance on the time it takes to make a response (Tullis, 2018); these cues likely will not be used in the same way by LLMs. Together, these differences in the assumed computations and inputs to metacognition may strongly impact how humans integrate LLMs’ expressed confidence into their own beliefs and decisions.
@@
Overall, it is clear that improving AI metacognition is a key priority: LLMs must be able to differentiate correct responses from incorrect ones. Yet our research trajectory must exceed simply improving LLMs’ self-evaluation capacities if they are to effectively collaborate with humans. Imbuing LLMs with appropriate metacognitive capacities must also include directed research into their communication of uncertainty to human users, and explicit comparisons between how humans and LLMs evaluate their own uncertainty. New tasks and evaluation strategies may be beneficial in driving such development, such as building LLM capacities to recognize and name skills required to solve the task at hand (e.g., mathematical problems) (Didolkar et al., 2024). Training regimes which drive alignment between LLMs’ verbalized confidence and the perceived confidence by humans (Stengel-Eskin et al., 2024), or which emphasize LLMs’ capacities to detect questions that are beyond the scope of their knowledge base or are unanswerable (Zhang et al., 2024), may also be powerful paths forward.
@@ 6 Future Benefits of Improved AI Metacognition
Beyond the importance of improving LLMs’ metacognitive capacities to facilitate their effective integration into human-AI joint decision-making, imbuing LLMs—or any AI system—with improved metacognition may also strongly support their approach to one ultimate target: generalized intelligence. In humans, metacognitive capacities facilitate goal-directed behaviors including learning, information-seeking, and more. For example, cognitive science has long recognized the role of metacognition in driving self-directed learning, which allows us to focus effort on acquiring information that we do not yet possess—with strong parallels to active learning AI algorithms that can optimally select their own training data to maximize efficient acquisition of coherent skills or beliefs (Gureckis & Markant, 2012). Metacognitive-like evaluations of uncertainty can drive information seeking behavior even in perception and learning, allowing agents to focus on acquiring knowledge to minimize uncertainty in their internal representations of the world (Schulz et al., 2023)—a form of curiosity-driven behavior. Confidence signals can also help agents learn in reinforcement learning contexts through explicit calculation of confidence-based prediction errors (Ptasczynski et al., 2022). Finally, meta-evaluations of one’s own metacognitive abilities can also drive humans’ learning (Recht et al., 2025), and the same could be true for AI systems. It is clear that promoting LLMs’ metacognitive capacities may have profound benefit for the development of smarter, more capable AI.

### v2

@@ 4.2 Divergences Between Human and LLM Metacognition
Despite a number of parallels, there remain important differences between human and LLM metacognition. In humans, many researchers suppose that the ability to form confidence judgments rests on the formation of a second-order representation: a separate evaluation or reassessment of the internal representations prompted by input information and which gave rise to a behavioral output (Peters, 2022). (Note: not all experts in human metacognition agree with the second-order assessment view; see Zheng et al. (2025) for discussion.) Unless explicitly present in their architecture, LLMs may not form such second-order self-evaluative representations unless explicitly prompted to do so. Relatedly, LLMs may be less able to correctly evaluate the source of uncertainty in their internal representations, suggesting they lag humans in distinguishing between metacognition and theory of mind. LLMs are prone to conflate their own beliefs with those attributed to others; that is, they are less able to separate the speaker’s belief from their own compared to humans when interpreting uncertain statements (Belém et al., 2024).
Another difference is the extent to which extent the metacognitive abilities can be improved through training. In the case of LLMs, research has shown that confidence verbalization can be improved through finetuning approaches that reward the LLM for accurately conveying uncertainty to a listener (Stengel-Eskin et al., 2024) or aligning overt confidence scores with implicit measures of uncertainty such as consistency scores (Steyvers et al., 2025a). Both metacognitive calibration and sensitivity are improved through training. However, the gains in metacognitive sensitivity tend to be domain specific and there is only limited generalizability to other knowledge domains and other types of questions (e.g., switching from multiple choice to short answers). For humans, providing providing feedback, encouraging reflective reasoning, and explicitly targeting cognitive biases can reduce human miscalibration of confidence (Kelly & Mandel, 2024; Rouy et al., 2022). However, there is no evidence that human metacognitive sensitivity improves in the presence of feedback (Haddara & Rahnev, 2022), likely reflecting underlying architectural differences: while LLMs’ metacognitive judgments can be fine-tuned through explicit training objectives, human metacognitive sensitivity appears to be constrained by more stable, possibly hardwired cognitive mechanisms that are less responsive to feedback.
Another difference may stem from the domain generality or specificity of metacognition in humans. It is thought that some shared processes underlie metacognition about perception, memory, and cognition may exist and rely on common neural structures, while others may be domain specific—i.e., separable computational or neural modules for perceptual versus cognitive or memory metacognition (Morales et al., 2018). A comprehensive assessment of the domain generality of LLMs’ metacognitive capacity has not yet been undertaken; however, preliminary evidence suggests that fine-tuning a model on a particular task (including training specific metacognitive capacities in that task) may not automatically generalize to other tasks (Steyvers et al., 2025a; Stengel-Eskin et al., 2024). As LLMs are increasingly integrated into many highly different tasks and reasoning domains, attending to their domain-specific versus domain-general metacognitive capacities will become increasingly urgent (see, e.g., (Griot et al., 2025) for LLMs’ metacognitive failures in medical reasoning).
@@ 5 Communication of Uncertainty in Human-AI Interaction
As discussed above, metacognitive sensitivity is the degree to which confidence judgments can discriminate between right and wrong answers, which is critical to effective decision-making in humans (Fleming, 2024). For optimal interaction and humans’ trust of AI systems, LLMs thus must be able to convey to human deciders whether their decisions are likely to be correct (Lee et al., 2025; Steyvers et al., 2025b; Li & Steyvers, 2025; Kadavath et al., 2022). Problematically, LLMs appear reluctant to express uncertainty (Zhou et al., 2024). Because humans rely heavily on linguistic uncertainty expressions (Steyvers et al., 2025b; Zhou et al., 2024), the absence of expressions of uncertainty may raise humans’ reliance on model outputs even beyond the already-overconfident judgments the models express. A potential reason for LLMs’ reluctance to express uncertainty may lie in the use of reinforcement learning from human feedback, where models are fine-tuned to produce outputs that align with human preferences. These preferences often favor responses that sound confident—even when that confidence may not reflect higher accuracy—leading LLMs to avoid verbal expressions of uncertainty during generation (Steyvers et al., 2025b; Zhou et al., 2024). Unfortunately, this problem may be further exacerbated as LLMs are used for increasingly challenging applications, potentially by increasingly non-expert users. Because individuals who do not possess topical expertise are less able to correctly assess the expertise of others (Bower et al., 2024), non-expert users may be especially influenced by superficial aspects of LLM responses—such as the absence of uncertainty expressions or the length of the answer. Recent findings show that users tend to interpret longer LLM responses as more confident, even when the model’s internal confidence remains unchanged (Steyvers et al., 2025b). This suggests that response length and style can mislead users into overestimating the certainty or reliability of the model’s output, potentially leading to overreliance on answers that do not warrant such confidence. Humans and LLMs may also rely on different sets of cues when assessing their confidence in other humans, such as humans’ reliance on the time it takes to make a response (Tullis, 2018); these cues likely will not be used in the same way by LLMs. Together, these differences in the assumed computations and inputs to metacognition may strongly impact how humans integrate LLMs’ expressed confidence into their own beliefs and decisions.
@@
Overall, it is clear that improving AI metacognition is a key priority: LLMs must be able to differentiate correct responses from incorrect ones. Yet our research trajectory must exceed simply improving LLMs’ self-evaluation capacities if they are to effectively collaborate with humans. Imbuing LLMs with appropriate metacognitive capacities must also include directed research into their communication of uncertainty to human users, and explicit comparisons between how humans and LLMs evaluate their own uncertainty. New tasks and evaluation strategies may be beneficial in driving such development, such as building LLM capacities to recognize and name skills required to solve the task at hand (e.g., mathematical problems) (Didolkar et al., 2024). Training regimes which drive alignment between LLMs’ verbalized confidence and the perceived confidence by humans (Stengel-Eskin et al., 2024), or which emphasize LLMs’ capacities to detect questions that are beyond the scope of their knowledge base or are unanswerable, may also be powerful paths forward.
@@ 6 Future Benefits of Improved AI Metacognition
Beyond the importance of improving LLMs’ metacognitive capacities to facilitate their effective integration into human-AI joint decision-making, imbuing LLMs—or any AI system—with improved metacognition may also play a role in progress toward more general forms of machine intelligence. In humans, metacognitive capacities—including metacognitive control, such as deciding what to learn and when—facilitate goal-directed behaviors including learning, information-seeking, and more. For example, cognitive science has long recognized the role of metacognition in driving self-directed learning, which allows us to focus effort on acquiring information that we do not yet possess (Gureckis & Markant, 2012). These curiosity-driven behaviors may reflect motivation to minimize uncertainty in our internal representations of the world (Schulz et al., 2023), with strong parallels to active learning AI algorithms that can optimally select their own training data to maximize efficient acquisition of coherent skills or beliefs (Gureckis & Markant, 2012). Confidence signals can also help agents learn in reinforcement learning contexts through explicit calculation of confidence-based prediction errors (Ptasczynski et al., 2022). Finally, meta-evaluations of one’s own metacognitive abilities can also drive humans’ learning (Recht et al., 2025), and the same could be true for AI systems. It is clear that promoting LLMs’ metacognitive capacities may significantly advance the design of AI systems with broader adaptive capacities.